OpenStudy (anonymous):

Jennifer wrote the following steps to simplify the expression: 3 – (x + 2x – 5).

OpenStudy (anonymous):

x = 8/3

OpenStudy (anonymous):

simplify the equation first.. 3-x-2x+5 8-3x equate to 0.. 8-3x=0 solve for x.. 8=3x x=8/3

OpenStudy (anonymous):

ummm.... i havent posted the full question yet.... just give me a moment...

OpenStudy (anonymous):

lol..

OpenStudy (anonymous):

3-x-2x+5 3-3x+5 3-3x=-5 -3x=-5-3=-8 x=8/3 negative and negative will become positive

OpenStudy (anonymous):

so mine ans is also correct

OpenStudy (anonymous):

Which step of justification is incorrect, and what should the justification for that step be to simplify the expression? Step 3; Associative Property of Addition Step 2; Associative Property of Addition Step 1; Commutative Property of Multiplication Step 4; Multiplication Property of Equality

OpenStudy (anonymous):

That's the full question. Sorry it took me a while to make the document, because it woulddnt let me do anything with the graphic.

OpenStudy (anonymous):

sorry....

OpenStudy (anonymous):

Sorry you wasted your time figuring out the wrong problem...